OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 12 and a13 = -60?

jimthompson5910 (jim_thompson5910):

an = a1 + d(n-1) a13 = 12 + d(13-1) -60 = 12 + d(13-1) solve for d and tell me what you get